let x be a random variable with probability density function given by f(x)=(2he^-1/4x ; 0<x find the value of constant h



1
Expert's answer
2022-01-26T18:25:38-0500
"\displaystyle\int_{-\infin}^{\infin}f(x)dx=1"

"\displaystyle\int_{0}^{\infin}2he^{-x/4}fdx=\lim\limits_{t\to\infin}\int_{0}^{A}2he^{-x/4}fdx"

"=\lim\limits_{t\to\infin}[-8he^{-x/4}]\begin{matrix}
 t\\
 0
\end{matrix}=8h=1"

"h=1\/8"